Osmotic pressure of a solution containing dissolved protein of solution is 20 MmHg at 27 degrees C calculate the concentration of protein


1
Expert's answer
2022-04-08T05:55:03-0400

"\pi" (solution) = 20 mmHg

T = 27 C = 300 K;

"\pi" (solution) = C * R * T;

R = 62.364 L * mmHg/(mol * K);

C = "\pi" (solution)/(R *T) = 20/(300 * 62.364) = 0.001 M.

Answer: 0.001 M.